# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v2

@@ 4.1. Mountain pass level estimate in the critical case
(5.20) $\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|$ $\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}\left(t_{\varepsilon}+\tfrac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{r-1}{\rm e}^{\alpha\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}|\widetilde{u}_{n}-\widetilde{u}|$
@@
$\lesssim\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{p-2}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x+\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{r}{\rm e}^{\alpha w_{n}^{2}}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|=:T_{1}+T_{2}\,,$
@@
(5.32) $S_{3}^{\overline{\nu}}$$\leq\int_{\{\widetilde{u}\leq\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}_{n}^{2}}\,{\rm d}x+\int_{\{\widetilde{u}>\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}^{2}}\,{\rm d}x\leq C$
@@
$\leq C+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}|^{pp^{\prime}}\,{\rm d}x\right)^{\frac{1}{p^{\prime}}}\!\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\,\widetilde{u}^{p}\,{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}pC_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{p}}\leq C\,,$
@@ Acknowledgements:
The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). Federico Bernini is partially supported by INdAM-GNAMPA Project 2024 titled Problemi spettrali e di ottimizzazione di forma: aspetti qualitativi e stime quantitative. Giulio Romani and Cristina Tarsi are partially supported by INdAM-GNAMPA Project 2023 titled Interplay between parabolic and elliptic PDEs and INdAM - GNAMPA Project 2024 titled New perspectives on Choquard equation through PDEs with local sources.

### v3

@@ 4.1. Mountain pass level estimate in the critical case
(5.20) $\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|$
$\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}\left(t_{\varepsilon}+\tfrac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{r-1}{\rm e}^{\alpha\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}|\widetilde{u}_{n}-\widetilde{u}|$
@@
$\lesssim\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{p-2}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
$\quad+\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{r}{\rm e}^{\alpha w_{n}^{2}}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|=:T_{1}+T_{2}\,,$
@@
(5.32) $S_{3}^{\overline{\nu}}$$\leq\int_{\{\widetilde{u}\leq\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}_{n}^{2}}\,{\rm d}x$
+$\quad+\int_{\{\widetilde{u}>\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}^{2}}\,{\rm d}x\leq C$
@@
$\leq C+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}|^{pp^{\prime}}\,{\rm d}x\right)^{\frac{1}{p^{\prime}}}\!\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\,\widetilde{u}^{p}\,{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}pC_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{p}}$
$\leq C\,,$
@@ Acknowledgements:
The main part of this work was carried out while F.B. was affiliated with Università degli Studi di Milano and G.R. with Università degli Studi dell’Insubria, whose support is gratefully acknowledged. The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). Federico Bernini is partially supported by INdAM-GNAMPA Project 2024 titled Problemi spettrali e di ottimizzazione di forma: aspetti qualitativi e stime quantitative. Giulio Romani and Cristina Tarsi are partially supported by INdAM-GNAMPA Project 2023 titled Interplay between parabolic and elliptic PDEs and INdAM - GNAMPA Project 2024 titled New perspectives on Choquard equation through PDEs with local sources.